OpenStudy (anonymous):

\sqrt[5]{256}div sqrt[6]{64} = 2^{x}

OpenStudy (anonymous):

\[\sqrt[5]{256}\div \sqrt[6]{64} = 2^{x}\]

OpenStudy (anonymous):

256 = 2^8 64 = 2^6 (2^8)^1/5 / (2^6) ^1/6 = 2^x 2^(8/5) / 2^1 = 2^x 2^(8/5 - 1) = 2^x 2^(3/5) = 2^x x= 3/5

OpenStudy (anonymous):

Crap, i forgot to drop the exponent of 1 from the denominator.... Thanks sector.

OpenStudy (anonymous):

by drop i mean bring it with for the rest of the party.